# English K_Quantization of LLMs Does Not Disproportionately Diminish Multilingual Performance

### v1

@@ Abstract
For consumer usage of locally deployed LLMs, the GGUF format and k_quantization are invaluable tools for maintaining the performance of the original model while reducing it to sizes deployable with consumer-grade hardware. The number of bits dedicated to each weight from the original model is reduced based on how important they are thought to be during model inference. This importance is arrived at through the application of an ’importance matrix’—a relatively small text document meant to be representative of the LLM’s standard use-cases. In the vast majority of quants available online, this document is primarily written in English. It was therefore an open question whether performance on English language tasks was preserved through the sacrifice of multilingual performance and whether it can be preserved with alternate importance matrices. This article investigates these hypotheses by quantizing Llama3.3 70B on importance matrices written in three different languages (English, Norwegian, and Malayalam) and evaluating them on the MixEval dataset in both English and Norwegian. All experiments related to k_quantization yielded non-significant results (In all cases p > 0.237) indicating that current quantization practices do not disproportionately harm multilingual performance.
@@ 1 Introduction
Large Language Models (LLMs) have had a remarkable impact on the daily lives of people over the last few years. Despite their global reach, these LLMs are generally formally trained to speak only a few widespread languages, such as English or Chinese. Considering the number of speakers for both English ($\sim 1.5B$), and Mandarin Chinese ($\sim 1.1B$), the choice to use these as a basis for an LLMs conversational ability is hardly surprising when compared to the $\sim 5.4M$ total Norwegian speakers[1]. However, these LLMs often acquire an ability to speak these relatively uncommon languages regardless through their inclusion in an LLM’s pre-training data[2]. For example, while non-English languages were described as out-of-scope for the original llama2 release, its pre-training dataset consisted of $10.3\%$ non-English text of which $0.03\%$ was Norwegian[3]. Until recently, only the extremely large proprietary LLMs of companies such as OpenAI or Anthropic could do so at a passable level, this capability was fully beyond the available open-source111including open weights models with a permissive license. Henceforth, these will both be referred to as ”open source” models available. This changed with the release of newer open-source models such as version Llama3.3[4] and Deepseek v3[5]. A qualitative sample of the difference between older and newer models is shown in Figure 4.
@@
The extreme multilingualism of modern LLMs has been achieved through many different training techniques, the most prominent of these being the sheer size of the models. The unique quirks of each language require the dedication of a significant number of weights, even if the exact nature and magnitude of these weights remains uncertain due to the black box nature of LLMs and Machine Learning (ML) as a whole [^2] We dive deeper into this topic in Section 2.1. The amount of memory required to store and run inference on the weights of these models often means that running them is prohibitively expensive for most consumers to run locally. In an effort to combat this problem, researchers have released methods for quantizing ML models, i.e. to reduce the number of bits per weight while retaining as much of the original behavior as possible. As the information known by an LLM is innately tied to its weights, this quantization process will naturally remove some content from its knowledge-base. It is uncertain, however, to what extent the multilinguality of a model is harmed by this process. In particular, in regards to a common method of quantizing .gguf files which employs the use of an importance matrix. This quantization method, and those similar to it, do not apply its quantization uniformly across all the weights of a model, rather it uses an example text document to assess the importance of each weight and dedicates fewer bits to those deemed of lesser importance to the functioning of the model. This importance metric text consists primarily of text written in English.
@@ 2 Related Work
A previous study by Marchisio et al. has investigated this topic for other quantization schemes and found the impact on multilingual performance to be negligible. Across several benchmarks, the performance drop for multilingual tasks is generally measured to be lower than 3%. This decline is more noticeable to human annotators, however, with an average performance drop by $13.1\%$ across all categories. Further, languages with a non-latin script are generally more severely affected by quantization on automated benchmarks[8].
@@ 2.1 Previous investigation into Multilingual LLMs
Some previous work exists investigating the inner workings of LLMs concerning multilinguality. Zhao et al. released a paper investigating the neurons of an LLM and their impact on its performance on multilingual tasks. They hypothesize that LLMs take multilingual input, translate it into English, process it, then translate it back into English.
@@ 3 Method
Few existing open source LLMs explicitly support Norwegian, however, most attain some level of proficiency in the language passively. We will be using llama3.3 as our initial experimental model, this family of models supports eight languages (English, German, French, Italian, Portuguese, Hindi, Spanish, and Thai)[4]. All experiments will be carried out on the relatively large llama3.3 70b LLMs due to their respectable performance on Norwegian language tasks. Based on experience, LLMs require roughly this number of parameters to perform well on such a minor language without having been explicitly trained to support it.
Several different datasets exist that may be used to evaluate the performance of an LLM, among the more established of these is MixEval, which was released in June 2024. The dataset notably has a 0.96 model rank correlation with the Chatbot Arena leaderboard[10]—one of the more well respected evaluation methods available today, without incurring the cost of human evaluation. As prominent datasets for LLM evaluation tend to contaminate training data, MixEval proposes a novel pipeline for dynamically building a dataset based on a mixture of existing LLM benchmarks. The dataset is updated regularly to maintain its integrity in the face of contamination. We will be using the $2024-06-01$ for our evaluation, a date well beyond the Llama3.3 data cut-off date of $2023-12-01$. The dataset is split into two portions, MixEval and MixEval-hard, in which the hard section represents questions that are deliberately picked to distinguish between stronger models. For this project we will be using the standard version of the dataset[11].
K-quant and the application of an importance matrix are methods for quantization developed by Iwan Kawrakow. While these methods have never been formally published in the form of a paper at the time of writing555”There are no papers on k- or i-quants because I don’t like writing papers. Combined with me enjoying the luxury of not needing another paper on my CV, and me not looking for a job or for investment, I see no reason to go and advertise on arXiv.” - Kawrakow[12], it has become ubiquitous due to its inclusion in the llamacpp library—as evidenced by how prevalent the quantization scheme is in for the .gguf format[13, 14]. While other approaches arrive at the importance matrix by looking at the weights gradient from a training run on a set of training tokens, this implementation calculates an approximation of these values based on the activations on a set of inferences.
@@ 3.1 Importance Matrix
One of the more prominent providers of llm gguf files is the Huggingface user Colin ’bartowski’ Kealty, who kindly provides the source text used for his importance matrices. The text consists of primarily English content describing several diverse domains compiled with the help of users "Dampf" and "Kalomaze". This source text has been machine-translated via gpt-4o into Norwegian for our own quantization purposes, and $0.63\%$ has been pruned due to being untranslatable. The pruned section primarily consisted of riddles built around spelling and wordplay. Further for our baseline comparison english importance matrix, the non-english sections have been removed.
@@ 4 Results & Discussion
Our overall findings indicate that English quantization is more performant than language-dedicated quantized models. With the starting hypothesis that quanting based on a non-English language significantly alters the performance on a benchmark, no p value is calculated to be lower than $0.2373$—a clear failure to disprove the null hypothesis. p values were calculated by Monte Carlo simulation with the assumption that the probability found by the English version of each quant tier models the underlying binomial distribution of the success chance for each question.
@@
It seems unlikely to us that the pipeline as mapped out by Zhao et al. in Figure 2 explains the entirety of multilingual behavior within LLMs. Firstly, because machine learning at this scale rarely results in such tidy behavior, and secondly, if this were the case, its ability to respond to multilingual queries would be equal to its ability to translate the given language. Qualitatively this is not the case, translation is a significantly simpler task than instruction following in our experience.
@@
The results for the free-form portion of MixEval can be seen in Figure 6. Similarly to the multiple choice questions, the accuracy of the different language quantization quant perform similarly to the English version.
@@
4.1 Future work
While the investigated quantization method deprioritizes weights that do not activate with the application of the importance matrix, it still attempts to retain their functionality in some form. The difference in quantization between languages detected by our methods was minimal, but an interesting alternate approach could be to apply a more deliberately destructive quantization method, for example, by scrambling the unused weights.
@@ 5 Conclusion
Our findings corroborate those of Marchisio et al.[8]. Quantization does not appear to disproportionately affect multilingual performance. Further, the usage of importance matrices written in non-English does not significantly improve performance on non-English datasets and might in fact slightly harm it. However, this reduction in performance is not statistically significant.

### v2

@@ Abstract
For consumer usage of locally deployed LLMs, the GGUF format and k_quantization are invaluable tools for maintaining the performance of the original model while reducing it to sizes deployable with consumer-grade hardware. The number of bits dedicated to each weight from the original model is reduced based on how important they are thought to be during model inference. This importance is arrived at through the application of an ’importance matrix’—a relatively small text document meant to be representative of the LLM’s standard use-cases. In the vast majority of quants available online, this document is primarily written in English. It was therefore an open question whether performance on English language tasks was preserved through the sacrifice of multilingual performance and whether it can be preserved with alternate importance matrices. This article investigates these hypotheses by quantizing Llama3.3 70B on importance matrices written in three languages (English, Norwegian, and Malayalam) and evaluating them on the MixEval dataset in both English and Norwegian. All experiments related to k_quantization yielded non-significant results (In all cases p > 0.237) indicating that current quantization practices do not disproportionately harm multilingual performance.
@@ 1 Introduction
Large Language Models (LLMs) have had a remarkable impact on the daily lives of people over the last few years. Despite their global reach, these LLMs are generally formally trained to speak only a few widespread languages, such as English or Chinese. Considering the number of speakers for both English ($\sim 1.5B$), and Mandarin Chinese ($\sim 1.1B$), the choice to use these as a basis for an LLMs conversational ability is hardly surprising when compared to the $\sim 5.4M$ total Norwegian speakers[1]. However, these LLMs often acquire an ability to speak these relatively uncommon languages regardless through their inclusion in an LLM’s pre-training data[2]. For example, while non-English languages were described as out-of-scope for the original llama2 release, its pre-training dataset consisted of $10.3\%$ non-English text of which $0.03\%$ was in Norwegian[3]. Until recently, only the extremely large proprietary LLMs of companies such as OpenAI or Anthropic could do so at a passable level, and this capability was fully beyond the available open-source111including open weights models with a permissive license. Henceforth, these will both be referred to as ”open-source” models available. This changed with the release of newer open-source models such as version Llama3.3[4] and Deepseek v3[5]. A qualitative sample of the difference between older and newer models is shown in Figure 4.
@@
The extreme multilingualism of modern LLMs has been achieved through many different training techniques, the most prominent of these being the sheer size of the models. The unique quirks of each language require the dedication of a significant number of weights, even if the exact nature and magnitude of these weights remains uncertain due to the black box nature of LLMs and Machine Learning (ML) as a whole [^2] We dive deeper into this topic in Section 2.1. The amount of memory required to store and run inference on the weights of these models often means that running them is prohibitively expensive for most consumers to run locally. In an effort to combat this problem, researchers have released methods for quantizing ML models, i.e. to reduce the number of bits per weight while retaining as much of the original behavior as possible. As the information known by an LLM is innately tied to its weights, this quantization process will naturally remove some content from its knowledge-base. It is uncertain, however, to what extent the multilinguality of a model is harmed by this process. In particular, in regards to the common methods of quantizing .gguf files that employs the use of an importance matrix, such as k_quantization. These quantization methods, and those similar to it, do not apply its quantization uniformly across all the weights of a model, rather it uses an example text document to assess the importance of each weight and dedicates fewer bits to those deemed of lesser importance to the functioning of the model. This importance metric text consists primarily of text written in English.
@@ 2 Related Work
A previous study by Marchisio et al. has investigated this topic for other quantization schemes and found the impact on multilingual performance to be present, but negligible. Across several benchmarks, the performance drop for multilingual tasks is generally measured to be lower than 3%. This decline is more noticeable to human annotators, however, with an average performance drop by $13.1\%$ across all categories. Further, languages with a non-latin script are generally more severely affected by quantization on automated benchmarks[8].
@@ 2.1 Previous investigation into Multilingual LLMs
Some previous work exists investigating the inner workings of LLMs concerning multilinguality. Zhao et al. released a paper looking into the neurons of an LLM and their impact on its performance on multilingual tasks. They hypothesize that LLMs take multilingual input, translate it into English, process it, then translate it back into English.
@@ 3 Method
Few existing open-source LLMs explicitly support Norwegian, however, most attain some level of proficiency in the language passively. We will be using llama3.3 as our initial experimental model. This family of models supports eight languages (English, German, French, Italian, Portuguese, Hindi, Spanish, and Thai)[4]. All experiments will be carried out on the relatively large llama3.3 70b LLMs due to their respectable performance on Norwegian language tasks. Based on experience, LLMs require roughly this number of parameters to perform well on such a minor language without having been explicitly trained to support it.
Several different datasets exist that may be used to evaluate the performance of an LLM. One of the more established of these is MixEval, which was released in June 2024. The dataset notably has a 0.96 model rank correlation with the Chatbot Arena leaderboard[10]—one of the more well respected evaluation methods available today, without incurring the cost of human evaluation. As prominent datasets for LLM evaluation tend to contaminate training data, MixEval proposes a novel pipeline for dynamically building a dataset based on a mixture of existing LLM benchmarks. The dataset is updated regularly to maintain its integrity in the face of contamination. We will be using the $2024-06-01$ for our evaluation, a date well beyond the Llama3.3 data cut-off date of $2023-12-01$. The dataset is split into two portions, MixEval and MixEval-hard, in which the hard section represents questions that are deliberately picked to distinguish between stronger models. For this project we will be using the standard difficulty version of the dataset[11].
K-quant and the application of an importance matrix are methods for quantization developed by Iwan Kawrakow. While these methods have never been formally published in the form of a paper at the time of writing555”There are no papers on k- or i-quants because I don’t like writing papers. Combined with me enjoying the luxury of not needing another paper on my CV, and me not looking for a job or for investment, I see no reason to go and advertise on arXiv.” - Kawrakow[12], it has become ubiquitous due to its inclusion in the llamacpp library—as evidenced by how prevalent the quantization scheme is for the .gguf format[13, 14]. While other approaches arrive at the importance matrix by looking at the weights gradient from a training run on a set of training tokens, this implementation calculates an approximation of these values based on the activations on a set of inferences.
@@ 3.1 Importance Matrix
One of the more prominent providers of llm gguf files is the Huggingface user Colin ’bartowski’ Kealty, who kindly provides the source text used for his importance matrices. The text consists of primarily English content describing several diverse domains compiled with the help of users "Dampf" and "Kalomaze". This source text has been machine-translated via gpt-4o into Norwegian for our own quantization purposes, and $0.63\%$ of it has been pruned due to being untranslatable. The pruned section primarily consisted of riddles built around spelling and wordplay. Further, for our baseline importance matrix in English, the non-English sections have been removed.
@@ 4 Results & Discussion
Our overall findings indicate that English quantization is more performant than language-dedicated quantized models. With the starting hypothesis that quanting based on a non-English language significantly alters the performance on a benchmark, no p value is calculated to be lower than $0.2373$ for the multiple choice dataset—a clear failure to disprove the null hypothesis. p values were calculated by Monte Carlo simulation with the assumption that the probability found by the English version of each quant tier models the underlying binomial distribution of the success chance for each question.
@@
It seems unlikely to us that the pipeline as mapped out by Zhao et al. in Figure 2 explains the entirety of multilingual behavior within LLMs. Firstly, because machine learning at this scale rarely results in such tidy behavior, and secondly, if this were the case, its ability to respond to multilingual queries would be equal to its ability to translate the given language. Qualitatively this is not the case, translation is a significantly simpler task than instruction following, in our experience. However, our results do not contradict their hypothesis as performance on both Norwegian and Malayalam tasks appear heavily connected to performance in English.
@@
The results for the free-form portion of MixEval can be seen in Figure 6. Similarly to the multiple choice questions, the accuracy of the different language quantization quant perform similarly to the English version. Both of the Norwegian and Malayalam quants of Q2_K_S arrive at p values lower than the standard significance treshold of 0.05 in English if these were evaluated in isolation. We should keep in mind that these experiments were performed with a common hypothesis for all of them, however, and therefore correct our significance treshold based on the number of experiments performed. By applying, for example, the Bonferroni correction[15] and adjusting the threshold to $\alpha=\frac{original\ \alpha}{numberofexperiments}$. The resulting significance threshold of $\frac{0.05}{12}=0.00417$ is still well below the lowest found p value.
@@
4.1 Limitations
Despite the number of experiments already performed, our investigations have only scratched the surface of the number of combinations of LLMs, languages, and quantization schemes to evaluate. A more thourough evaluation would account for a wider range of these combinations, however, the results in this article should be good enough to given an indication of the rest of the field.
While the investigated quantization method deprioritizes weights that do not activate with the application of the importance matrix, it still attempts to retain their functionality in some form. The difference in quantization between languages detected by our methods was minimal, but an interesting alternate approach could be to apply a more deliberately destructive quantization method, for example, by deactivating the unused weights.
@@ 5 Conclusion
Our findings corroborate those of Marchisio et al.[8]. Quantization does not appear to meaningfully disproportionately affect multilingual performance. Further, the usage of importance matrices written in non-English does not significantly improve performance on non-English datasets and might in fact slightly harm it. However, this reduction in performance is not statistically significant.